I can not figure out how to put this in algebraic form to work it out, here is the problem...........Amy is choosing a cell phone plan. Three different companies offer a different number of free minutes of phone calls per month. City Com offers 280 less than twice the number of free minutes offered by Talk for Less Phone. Renee's Cell Phone offers 80 more free minutes per month than Talk for Less Phone. The sum of the free minutes offered by City Com and Talk for Less Phone is equal to twice the number of free minutes offered by Renee's Cell Phone. How many free minutes does each company offer?
Answer by ankor@dixie-net.com(22740)

Three different companies offer a different number of free minutes of phone calls per month.
City Com offers 280 less than twice the number of free minutes offered by Talk for Less Phone.
Renee's Cell Phone offers 80 more free minutes per month than Talk for Less Phone.
The sum of the free minutes offered by City Com and Talk for Less Phone is
equal to twice the number of free minutes offered by Renee's Cell Phone.
How many free minutes does each company offer?
:
Let m = no. of free minutes offered by "Talk for Less"
:
Write an equation for each statement:
:
" City Com offers 280 less than twice the number of free minutes offered by Talk for Less Phone."
2m - 280 = City Com's free minutes
:
"Renee's Cell Phone offers 80 more free minutes per month than Talk for Less Phone."
m + 80 = Renee's free minutes
:
"The sum of the free minutes offered by City Com and Talk for Less Phone is
equal to twice the number of free minutes offered by Renee's Cell Phone."
(2m-280) + m = 2(m+80)
2m - 280 + m = 2m + 160
3m - 2m = 160 + 280
m = 440 min offered by t. for less
:
you should be able to find the no. of minutes of the other companies
Check the solutions in the last statement